MIDC News
Year of 2016
December 29, 2016
- 08:26 MST: The SolarTAC Trimble (for PWV) was relocated on December 27 from one of the SolarTAC buildings to the instrument platform with the solar tracker. PWV data on the web site will not update until UCAR can do a new coordinate solation for the Trimble. -Afshin Andreas
- 09:28 MST: The SRRL BMS HMP-60 Temp/RH Probe (Deck) was swapped out with a calibrated sensor at 09:23 MST today. -Afshin Andreas
- 15:59 MST: The SRRL BMS HMP-60 Temp/RH Probe (Tower) was swapped out with a calibrated sensor at 15:36 MST today. -Afshin Andreas
- 09:57 MST: SRRL BMS Snow Depth sensor failed on December 6. Problem unknown at this time. -Afshin Andreas
- 13:49 MST: November 2016 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 13:50 MST: A new Global Normal LI-200R was added to SRRL BMS on Dec 1 @ 13:47 MST. The pyranometer is mounted next to the Global Normal CMP11 on the radiometer tower. -Afshin Andreas
- 10:41 MST: As of 10:05 MST today, the SRRL PGS-100 spectroradiometer is now collected by a CR800 data logger for higher reliability, onsite data storage, and consistency with the EKO spectroradiometers. -Afshin Andreas
- 08:47 MST: As of 14:20 MST yesterday, the OTF MS-700 spectroradiometer is now collected by a CR800 data logger at the OTF for higher reliability, onsite data storage, and consistency with the SRRL MS-7XX. -Afshin Andreas
- 09:50 MST: October 2016 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 16:23 MST: The SRRL SolarSIM-D2+ was calibrated on November 3, 2016. The new calibration is effective for all data collected by the instrument since it was installed (September 1, 2016). -Afshin Andreas
- 12:41 MST: The OTF MS-700 was calibrated against NIST FEL Standard of Spectral Irradiance S/N F680 on 9-NOV-2016 and returned to service on 10-NOV-2016 at 16:20 MST. The instrument was received out of tolerance with a 7% loss in response (all wavelengths). -Afshin Andreas
- 14:22 MST: The OTF MS-700 spectroradiometer was removed from service at 11:34 MST for calibration. -Afshin Andreas
- 08:01 MST: The SRRL MS-7XX Spectroradiometer was returned to back to 5-minute scans at 7:50 MST today. -Afshin Andreas
- 07:55 MST: The SRRL MS-7XX spectroradiometer shutdown yesterday @ 15:28 MST and was returned service today @ 7:20 MST in global horizontal orientation. -Afshin Andreas
- 11:10 MST: The SRRL BMS Direct Normal LI-190 PAR sensor was calibrated yesterday. New factor will be entered into logger at midnight. -Afshin Andreas
- 11:07 MST: Today @ 10:44 MST the scan rate of the MS-7XX spectroradiometer was changed from 5-minute to 2-minute. It will stay that way for the duration of the week. -Afshin Andreas
- 13:43 MST: The SRRL BMS POA LI-200 sensor (on SunEdison PV string) was replaced today at 13:40 MST with a sensor calibrated in BORCAL 2016-02. Logger program was updated at 13:42 MST. -Afshin Andreas
- 16:17 MST: The on-the-fly spectral data downloads for the SolarSIM have been updated to include 3 additional columns vs wavelength from the model, this includes: AOD, Ozone Transmission, and Water Vapor Transmission. -Afshin Andreas
- 11:19 MST: The shading ball has been removed from the diffuse radiometer at UTPA SRL on Oct 20 @ 14:00 MST. Data will read global until maintenance has been completed. -Afshin Andreas
- 13:18 MST: The SRRL BMS TWC Photometer was calibrated today (the sensor was misaligned during April 2016 calibration). The Vo for the TWC was updated at 13:16 MST. -Afshin Andreas
- 14:39 MST: A SolarSIM-D2 was installed at SRRL on September 1, 2016. Raw data and modeled spectral data can be generated on-the-fly by the MIDC web site. -Afshin Andreas
- 13:42 MST: September 2016 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 12:58 MST: At @ 12:40 MST the scan rate of the MS-7XX spectroradiometer was changed back to 5-minutes and the logger was changed back to a CR1000. Note: It takes 2 minutes for the CR1000 to process the data. -Afshin Andreas
- 13:38 MST: The SRRL BMS Direct RG780 NIP has been calibrated. New calibration factor will be applied on 10/1/2016 at 0:00 MST. -Afshin Andreas
- 12:34 MST: August 2016 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 09:44 MST: The DNI and DHI sensors at SolarTAC were swapped with calibrated sensors between 9:14 and 9:36 MST today. Data logger program was updated with new factors at 9:42 MST. -Afshin Andreas
- 16:23 MST: On Sept 20 @ 10:59 MST the scan rate of the MS-7XX spectroradiometer was changed from 5-minute to 1-minute. It will stay that way for the duration of the NREL Pyrheliometer Comparison (NPC). This required temporarily changing the datalogger from a CR1000 to a faster CR3000. -Afshin Andreas
- 15:34 MST: The SRRL MS-7XX spectroradiometer system was returned to service at 15:30 MST today. Instrument was calibrated against NIST FEL Standard of Spectral Irradiance S/N F680 and will be measuring direct normal until further notice. -Afshin Andreas
- 14:04 MST: The SRRL Global Horizontal MS-7XX spectroradiometer was removed from service at 13:45 MST today. It will be re-calibrated and then re-oriented for direct normal measurements later this week. -Afshin Andreas
- 14:14 MST: The University of Oregon Solar Radiation Monitoring Laboratory (SRML) has been added to MIDC. Existing PWV data for UofO has been moved to the same SRML station page. -Afshin Andreas
- 12:04 MST: Communications has been restored to NWTC M2. Calibration constants for new temperature and wind speed sensors were updated in logger program at 9:28 MST today. -Afshin Andreas
- 16:49 MST: On August 1, 2016 @ 13:45 MST, a Global Horizontal Hukseflux SR25 pyranometer was added to SRRL BMS. This instrument has a sapphire dome and currently no corrections are done on the data. -Afshin Andreas
- 13:47 MST: The NWTC M2 tower was dropped on 8/12/2016 at 07:13 MST, so that the temperature probes, wind speed and direction sensors could replaced. The tower was returned to service at approx. 10:20 MST, however the wireless Ethernet had failed, estimated repair date by IT department is 8/16. -Afshin Andreas
- 13:38 MST: July 2016 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 10:38 MST: The SRRL BMS Pressure sensor was swapped out with a calibrated sensor @ 09:55 MST today. -Afshin Andreas
- 13:18 MST: The following SRRL BMS Radiometer Tower instruments were swapped out today between 12:20 and 13:00 MST with sensors calibrated in BORCAL SW 2016-02: Upwelling CM21, CM3 & LI-200; Global 40-South CMP11 & LI-200; Global Normal CMP11; Global 90 North/East/South/West PSP & LI-200. Calibration factors were updated in logger at 12:25 MST for Upwelling CM21; 12:31 MST for remaining non-90 degree tilt; 12:52 MST for (8) 90 degree tilts. -Afshin Andreas
- 15:23 MST: The following SRRL BMS Diffuse instruments were swapped out with instruments that were calibrated in BORCAL 2016-03 between 15:10 to 15:20 MST today: CM22 Secondary. Cal factors were updated in the logger @ 15:20 MST today. No data loss occurred for CM22 Secondary (because data was automatically replaced with data from the CM22 Primary). -Afshin Andreas
- 12:03 MST: The SRRL BMS Direct TUVR, CUVA2, and CUVB2 were calibrated today. Data logger was updated with new calibration factors @ 12:01 MST. -Afshin Andreas
- 13:54 MST: June 2016 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 15:37 MST: A Black Photon BPI-IT1 direct normal instrument with top, middle, and bottom cells was installed at SolarTAC on June 9, 2016. -Afshin Andreas
- 12:37 MST: The SRRL LI-1800 Global 1-Axis spectroradiometer was returned to service at 12:30 MST today. A/D board was replaced with a spare and instrument was recalibrated. -Afshin Andreas
- 13:36 MST: The SRRL BMS TWC Photometer and Global RG780 PSP have been calibrated. The Vo for the TWC was updated at 13:31 MST and the CF for the RG780 was updated at 13:35 MST. -Afshin Andreas
- 06:12 MST: Calibration factors for the SRRL BMS Global PQS1 and LI-210 quantum sensors, and the LI-190 photometeric sensor was updated at 6:09 MST. The Upwelling LI-210 quantum sensor was updated at 6:08 MST. -Afshin Andreas
- 11:19 MST: The option to sum spectral data in the UV-Vis (UV-Visible) range has been added to MIDC. This feature will be used to help calibrate BMS RG780 radiometers. -Afshin Andreas
- 15:14 MST: The following SRRL BMS Diffuse instruments were swapped out with instruments that were calibrated in BORCAL 2016-02 between 14:28 to 14:46 MST today: CM22 Primary, 8-48, and CMP11. Cal factors were updated in the logger @ 14:47 MST today. No data loss occurred for CM22 Primary and 8-48 (because data was automatically replaced with data from the CM22 Secondary). -Afshin Andreas
- 13:47 MST: May 2016 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 15:33 MST: SRRL BMS deck data logger program was updated at 14:58 MST today with new calibration factors for instruments that were calibrated in BORCAL 2016-01, this includes Globals: CMP22 (vent), CMP22 #1, CMP22 #2, CMP11, CM6b, CM3, MS802, MS410, MS602, SPP, GPP, TSP-700, SPN1, SPLite2, ML-01, LI-200R, LI-200, SP-110; Directs: CHP1 #1, CHP1 #2, sNIP, NIP, DR02, MS-56, LI-201. Also the 2-degree responsivities for the PSP were updated at 15:00 MST today. -Afshin Andreas
- 09:28 MST: January 2015 through April 2016 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 07:45 MST: The NWTC M2 Global PSP was swapped out with a calibrated sensor at 6:45 MST today. Calibration factor was updated in data logger at 7:40 MST today. -Afshin Andreas
- 15:17 MST: Power to SolarTAC tracker was restored at 13:17 MST today. -Afshin Andreas
- 09:49 MST: Power to SolarTAC tracker was shut off at 09:35 MST today for electrical disconnect installation. Direct and Diffuse will be bogus during this time. -Afshin Andreas
- 08:16 MST: The SRRL LI-1800 Global 1-Axis spectroradiometer failed at sunrise on March 23. It appears to be a mechanical problem. -Afshin Andreas
- 08:14 MST: Due to a scheduled SRRL power outage, to replace the data lab humidifier, the following BMS instruments were down on March 29 from 7:40 to 7:54 MST: TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100, and SR50A. -Afshin Andreas
- 16:31 MST: The SRRL PGS-100 spectroradiometer was calibrated on March 10, 2016 outdoors against the NIST traceable LI-1800 PRS-175 spectroradiometer. The new calibration file was applied to data starting on 3/17/2016 at 16:40 MST. -Afshin Andreas
News Archives:
[1997] [1998] [1999] [2000] [2001] [2002] [2003] [2004] [2005] [2006] [2007] [2008] [2009] [2010] [2011] [2012] [2013] [2014] [2015] [2016] [2017] [2018] [2019] [2020] [2021] [2022] [2023]